OpenStudy (anonymous):

identify the vertex and the axis of symmetry for the graph of y=-5(x-2)^2+4

sam (.sam.):

Vertex (2,4) Axis of symmetry =2

OpenStudy (anonymous):

how did you find the axis of symmetry?

sam (.sam.):

From your equation y=-5(x-2)^2+4 (x-2)=0 x=2

sam (.sam.):

The vertex form of a quadratic is given by y = a(x – h)2 + k, where (h, k) is the vertex.

OpenStudy (anonymous):

okay thankyou :)

sam (.sam.):

yw

OpenStudy (asnaseer):

You might find this site useful if you want to understand this further: http://www.mathwarehouse.com/geometry/parabola/standard-and-vertex-form.php